# Gauge-independent gravitational waves from a minimal dark $U(1)$ sector with viable dark matter candidates

### v1

@@ 3.2.2 Phase transition at zero temperature
where $Z(\phi_{c})$ denotes the field renormalization factor and $C_{0},D,\tilde{D}$ are Nielsen coefficients; the total derivative term containing $\tilde{D}$ are discussed in [39]. Substituting the derivative expansions given in Eqs. (3.26)–(3.28) into Eq. (3.24), the Nielsen identity at order $\partial^{0}$ and $\partial^{2}$ is derived as
@@
in contrast to the usual Higgs potential with a negative mass-squared term. Additionally, since the barrier is generated by the Coleman-Weinberg contribution, the positive mass term $\sim\mu^{2}$ cannot be parametrically large, otherwise the potential will not develop. This setup can thus support a zero-temperature phase transition and is often adopted for pioneering studies [37].
@@
$=\frac{g_{x}^{2}}{16\pi^{2}}\Bigl(\xi\log\frac{m_{c}^{2}}{\Lambda^{2}}+3\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}+\xi\Bigr)\,$, (3.42)
and perturbation analysis shows that the Nielson coefficient $C_{0}(\phi_{c})$ starts at order $g_{x}^{2}$ and $D,\tilde{D}$ starts at unity, expressed as
@@
The standard procedure for calculating the bubble nucleation rate at zero temperature is finding a bounce solution $\phi_{b}$ [78, 79] from the leading-order equation of motion
$\Box\phi_{b}=\frac{\partial V^{0{\rm T}}_{g_{x}^{4}}}{\partial\phi_{c}}\biggr|_{\phi_{c}=\phi_{b}}\,.$ (3.48)
@@ 3.2.3 Phase transition at high temperatures ($T\gg m_{A^{\prime}}$)
The negative mass term $\mu_{x}^{2}$ competes with the positive $T^{2}$ corrections, leading to a critical crossover temperature $T_{o}$ where the Higgs effective mass is zero, i.e., $\mu_{{\rm eff}}^{2}=0$. This temperatures is just slightly below $T_{c}$ and a little bit above $T_{n}$, i.e., $T_{c}\gtrsim T_{o}\gtrsim T_{n}$. Near the high-temperature phase transition, all terms in the potential are expected to be of the same order of magnitude [3, 42], [^5] The full thermal potential can be effectively written as $V(\phi,T)=D(T^{2}-T_{0}^{2})\phi^{2}-ET\phi^{3}+\frac{\lambda_{x}}{4}\phi^{4}\,$, (3.57) where $T_{0}$ is the temperature at which $\phi=0$ is no longer a local minimum. At the critical temperature $T_{c}$, $V^{\prime}(\phi,T_{c})=0\,,\qquad V(\phi,T_{c})=V(0,T_{c})\,$, (3.58) which solve to $\phi_{c}=\frac{2E}{\lambda_{x}}T_{c}\,,\qquad D(T^{2}-T_{0}^{2})=\frac{E^{2}}{\lambda_{x}}T_{c}^{2}\,,\qquad{\rm at}\ T_{c}\,.$ (3.59) Plugging in back to the potential Eq. (3.57) we find all three terms are the same parametric size $\sim E^{4}T_{c}^{4}/\lambda_{x}^{3}$. At $T_{n}$ which lies just slightly below $T_{c}$, relation Eq. (3.59) approximately holds and thus $\frac{\phi_{c}}{T_{n}}\approx\frac{2E}{\lambda_{x}}\sim\mathcal{O}(1)\,.$ (3.60) In gauge theory, $E\sim g_{x}^{3}$ and thus $\lambda_{x}\sim g_{x}^{3}$. yielding the following power counting relation:
@@
Using the power counting introduced in Eq. (3.61), the leading-order effective potential is obtained as
@@
$(m_{G}^{2}+\Pi_{G})^{3/2}-m_{c}^{3}$ $=\left(-\mu_{{\rm eff}}^{2}+\lambda_{x}\phi_{c}^{2}+\xi g_{x}^{2}\phi_{c}^{2}+\frac{1}{4}\lambda_{x}T^{2}\right)^{3/2}-\left(\xi g_{x}^{2}\phi_{c}^{2}\right)^{3/2}$
$\approx\left(-\mu_{{\rm eff}}^{2}+\lambda_{x}\phi_{c}^{2}+\xi g_{x}^{2}\phi_{c}^{2}\right)^{3/2}-\left(\xi g_{x}^{2}\phi_{c}^{2}\right)^{3/2}\sim\sqrt{\xi}\,\mathcal{O}(g_{x}^{4})\,$, (3.71)
@@
For nucleation-related observables, one typically requires higher accuracy than that provided by the leading-order barrier alone. The next-to-leading order (NLO) effective potential is derived within a three-dimensional finite-temperature effective field theory (EFT) framework. The three-dimensional effective potential is computed in a constant background field $\phi_{3}$ using three-dimensional perturbation theory: the leading-order piece is the tree-level potential plus the leading three-dimensional one-loop contributions, while the NLO piece consists of the relevant two-loop vacuum diagrams in the three-dimensional EFT together with the corrections induced by NLO-matched parameters. After renormalization and consistent power counting, one arrives at [43, 44]
@@
The leading-order potential $V_{g_{x}^{3}}$, c.f., Eq. (3.66), is manifestly gauge-independent, and the corresponding bounce solution $\phi_{b}$ satisfies
$\nabla^{2}\phi_{b}=\frac{\partial V_{g_{x}^{3}}}{\partial\phi_{c}}\biggr|_{\phi_{c}=\phi_{b}}\,,\qquad\phi_{b}(\infty)=0\,,\qquad\phi_{b}^{\prime}(0)=0\,.$ (3.79)
@@
The leading order effective action is gauge-independent, since the leading order effective potential $V_{g_{x}^{3}}$ is constructed to be $\xi$-independent.
@@ 3.2.4 Phase transition at low temperatures ($T\ll m_{A^{\prime}}$)
Unlike the zero-temperature toy setup with a positive Higgs mass term, a minimal gauged $U(1)$ Higgs sector contains the standard tachyonic mass-squared term. In the supercooled regime, thermal effects, although being small since $T_{p}\ll v_{x}$, can lift the effective mass term positive and thereby create and maintain a false vacuum at $\phi=0$, as long as the tachyonic mass-squared parameter $\mu_{x}^{2}$ is parametrically small.
@@
Therefore, away from the barrier region, thermal effects are generally negligible because $T\ll\phi_{c}(T)$, and all thermal contributions can be safely neglected relative to the purely field-dependent terms.
Nevertheless, thermal effects are essential for generating the barrier around $\phi\sim T_{p}/g_{x}$, enabling the supercooled phase transition. In the vicinity of $\phi\sim T_{p}/g_{x}$, the temperature-dependent contributions can be comparable to the field-dependent terms. We therefore retain all thermal contributions in both the leading and subleading orders of effective potentials. Consequently, the leading and subleading order effective potentials are thus written as
@@
$\quad+V_{{\rm daisy}}^{A^{\prime}}(T^{3})+V_{{\rm daisy}}^{h_{x}}(T^{3})+\frac{T^{4}}{2\pi^{2}}\left[J_{B}\Bigl(\frac{m_{h_{x}}^{2}}{T^{2}}\Bigl)+3J_{B}\Bigl(\frac{m_{A^{\prime}}^{2}}{T^{2}}\Bigl)\right]\,$, (3.93)
@@
$\quad+V_{{\rm daisy}}^{G}(T^{3})+\frac{T^{4}}{2\pi^{2}}\left[J_{B}\Bigl(\frac{m_{G}^{2}}{T^{2}}\Bigr)-J_{B}\Bigl(\frac{m_{c}^{2}}{T^{2}}\Bigr)\right]\,$, (3.94)
@@
where we incorporate the gauge-independent thermal contributions into the leading order effective potential, while the $\xi$-dependent thermal terms are assigned to the subleading effective potential.
@@ 3.2.5 Comments and discussions
  • The power counting rule: For the high-temperature phase transition analysis in [39, 43, 44], the Higgs quartic coupling is taken to scale as $\lambda_{x}\sim g_{x}^{3}$. More precisely, it is assumed to lie in the region $g_{x}^{4}\ll\lambda_{x}\ll g_{x}^{3}$ [3]. [^6] For high-temperature phase transitions, one expects $\frac{g_{x}\phi_{c}}{T_{n}}\sim\frac{g_{x}\phi_{c}}{T_{c}}\ll 1\,.$ (3.104) Using the thermal effective potential in Eq. (3.57), one finds at $T_{c}$, $\frac{\phi_{c}}{T_{c}}\sim\frac{2E}{\lambda_{x}}\sim\frac{g_{x}^{3}}{\lambda_{x}}\,.$ (3.105) Substituting this into the previous condition gives $g_{x}\,\frac{\phi_{c}}{T_{c}}\sim\frac{g_{x}^{4}}{\lambda_{x}}\ll 1\,.$ (3.106) Hence, a consistent high-temperature phase transition typically requires $\lambda_{x}\gg g_{x}^{4}$, implying that, for fixed $\lambda_{x}$, moving deeper into the high-temperature regime favors smaller values of $g_{x}$. In our parameter scan, we find that all viable high-temperature phase transitions satisfy $\lambda_{x}\lesssim g_{x}^{3}$, while in many cases $\lambda_{x}\ll g_{x}^{4}$. This typically occurs for benchmark points in which $T_{n}$ lies somewhat below $T_{c}$, while the high-temperature approximation $m_{A^{\prime}}/T_{n}\ll 1$ remains reasonably valid.
    Nevertheless, adopting the scaling $\lambda_{x}\sim g_{x}^{3}$ to organize the leading and subleading terms in the potential does yield a gauge-independent effective action, as can be verified analytically using the Nielsen identity. However, alternative scalings, e.g., $\lambda_{x}\sim g_{x}^{4}$, do not yield a consistent result, as the Nielsen identity is not satisfied at the corresponding order, and thus the resulting subleading effective action is not gauge-independent.
@@
  • Gauge-parameter dependence: In the gauge-independent treatment, one must ensure that the terms neglected from the full effective potential are indeed parametrically small. Since some of the omitted contributions scale as $\sim\xi^{n}g^{\text{higher powers}}$, the gauge parameter $\xi$ cannot be taken arbitrarily large. Otherwise the power counting scheme ceases to be reliable and the expansion breaks down, even if the Nielsen identity is satisfied up to subleading order in the effective potential. In practice, we find that the gauge-independent treatment remains valid for moderately small values of $\xi$, e.g., $\xi\lesssim\mathcal{O}(1-10)$.
@@
  • Comments on the low-temperature analysis: Our scan indicates that a supercooled phase transition typically favors a moderately large gauge coupling, [^7] A conformal $U(1)$ Higgs sector typically contains no tree-level $\phi^{2}$ term. The local minimum at the origin is instead stabilized by the finite-temperature thermal mass. In this setup, a potential barrier can arise even without relying on the thermal cubic contribution $\sim T\,\phi^{3}$, and may be generated predominantly by the Coleman-Weinberg potential. As a result, the allowed range of the gauge coupling $g_{x}$ can be more flexible than in the conventional Higgsed $U(1)$ case with a tachyonic mass term. Moreover, conformal $U(1)$ models naturally admit shallow potentials, and it is therefore common to find $T_{n,p}\ll v$ and $\alpha\gtrsim 1$, corresponding to strong supercooled phase transitions that yield enhanced gravitational wave signals with peak frequencies shifted to lower bands [10, 52, 54, 80, 21, 31, 63]. Hence conformal hidden sectors are often discussed as natural pathways to nanohertz band signals. e.g., $g_{x}\gtrsim 0.5$, in sharp contrast to the high-temperature case, where smaller $g_{x}$ is generally preferred, c.f, Eq. (3.106). This trend can be understood from two main considerations.
    First, as discussed in the previous section, the low-temperature barrier is generated by the thermal cubic term $\sim T\phi^{3}$ near $\phi=0$, whose coefficient scales as $E\propto g_{x}^{3}$. A larger $g_{x}$ therefore enhances the cubic term and makes it easier to develop a sufficiently strong barrier for the supercooled phase transition. [^8] The effective thermal potential in Eq. (3.57) is derived in the high-temperature limit, where the $J_{B}$ functions admit a high-temperature expansion. In the low-temperature regime, Eq. (3.57) remains applicable in the field range $\phi\lesssim T_{p}/g_{x}$. In particular, as $\phi\to 0$ the field value is again much smaller than the temperature, and thus the high-temperature expansion remains valid in this small-field region, which is the most relevant for generating a thermal barrier. Nevertheless, Eq. (3.57) should not be used indiscriminately across the full temperature range in analyses of supercooled phase transitions.
    Second, the tree-level tachyonic mass term must be lifted to an effective positive value by thermal mass corrections proportional to $g_{x}^{2}$ in order to create and stabilize a local minimum at $\phi=0$. A larger $g_{x}$ thus keeps the symmetric point $\phi=0$ metastable down to lower temperatures, delaying the loss of metastability, which is essential for realizing a supercooled transition.

### v2

@@ 3.2.1 The field-dependent effective action
Bubble nucleation is governed by the Euclidean effective action evaluated on the bounce configuration. In practice, $\Gamma_{E}$ is approximated by a derivative expansion truncated at $\mathcal{O}(\partial^{2})$: at zero temperature this yields Eq. (3.49); while at finite temperature it leads to Eqs. (4.3) and (4.4). Accordingly, the Euclidean actions $S_{E}$ employed later in the paper are controlled approximations to the underlying 11PI effective action $\Gamma[\phi]$, evaluated on the relevant bounce configuration $\phi_{b}$.
@@ 3.2.2 Phase transition at zero temperature
where $Z(\phi_{c})$ denotes the field renormalization factor and $C_{0},D,\tilde{D}$ are Nielsen coefficients; the total derivative term containing $\tilde{D}$ is discussed in [43]. Substituting the derivative expansions given in Eqs. (3.27)–(3.29) into Eq. (3.25), the Nielsen identity at $\mathcal{O}(\partial^{0})$ and $\mathcal{O}(\partial^{2})$ is derived as
@@
in contrast to the usual Higgs potential with a negative mass-squared term. Additionally, since the barrier is generated by the Coleman-Weinberg contribution, the positive mass term $\sim\mu^{2}$ cannot be parametrically large, otherwise the barrier will not develop. This setup can thus support a zero-temperature phase transition and is adopted for pioneering studies [41].
@@
$=1+\frac{g_{x}^{2}}{16\pi^{2}}\Bigl(\xi\log\frac{m_{c}^{2}}{\Lambda^{2}}+3\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}+\xi\Bigr)+\mathcal{O}(g_{x}^{4})\,$, (3.43)
and perturbation analysis shows that the Nielsen coefficient $C_{0}(\phi_{c})$ starts at order $g_{x}^{2}$ and $D,\tilde{D}$ starts at unity, expressed as
@@
At zero temperature, the tunneling rate is governed by the four-dimensional Euclidean action
$S_{E}(\phi)=\int{\rm d}^{4}x_{E}\left[\frac{1}{2}\,(\partial_{\mu}\phi)^{2}+V(\phi)\right]\,,\qquad x_{E}=(\tau,\vec{x})\,.$ (3.49)
Varying $S_{E}$ with respect to $\phi$ leads to a partial differential equation for the bounce configuration. The solution that extremizes $S_{E}$ is invariant under rotations in Euclidean space and hence possesses an $O(4)$ symmetry. It thus depends only on the radial coordinate $\rho=\sqrt{\tau^{2}+x^{2}+y^{2}+z^{2}}$, i.e., $\phi=\phi(\rho)$. The field equation then reduces to the ordinary differential equation
$\phi^{\prime\prime}(\rho)+\frac{3}{\rho}\,\phi^{\prime}(\rho)=\frac{{\rm d}V}{{\rm d}\phi}\,$, (3.50)
subject to boundary conditions
$\phi\,(\rho\to\infty)=0,\qquad\left.\frac{{\rm d}\phi}{{\rm d}\rho}\right|_{\rho=0}=0\,.$ (3.51)
The gauge-independent procedure for computing the zero-temperature bubble nucleation rate proceeds by determining the bounce solution $\phi_{b}$ [82, 83] from the leading-order effective potential, which is automatically independent of the gauge-fixing parameter $\xi$
$\Box\phi_{b}=\frac{\partial V^{0{\rm T}}_{g_{x}^{4}}}{\partial\phi_{c}}\biggr|_{\phi_{c}=\phi_{b}}\,,\qquad\phi_{b}(\infty)=0\,,\qquad\phi_{b}^{\prime}(0)=0\,.$ (3.52)
@@ 3.2.3 Phase transition at high temperatures ($T\gg m_{A^{\prime}}$)
The negative mass term $\mu_{x}^{2}$ competes with the positive $T^{2}$ corrections, leading to a critical crossover temperature $T_{o}$ where the Higgs effective mass is zero, i.e., $\mu_{{\rm eff}}^{2}=0$. This temperature is just slightly below $T_{c}$ and a little bit above $T_{n}$, i.e., $T_{c}\gtrsim T_{o}\gtrsim T_{n}$. Near the high-temperature phase transition, all terms in the potential are expected to be of the same order of magnitude [3, 46], [^5] Using the high-temperature effective thermal potential given in Eq. (3.16), the critical temperature $T_{c}$ is defined by $V^{\prime}(\phi,T_{c})=0\,,\qquad V(\phi_{c},T_{c})=V(0,T_{c})\,$, (3.61) where $\phi_{c}\neq 0$ denotes the broken-phase minimum at $T_{c}$. Solving these conditions yields $\phi_{c}=\frac{2E}{\lambda_{x}}T_{c}\,,\qquad D(T^{2}-T_{0}^{2})=\frac{E^{2}}{\lambda_{x}}T_{c}^{2}\,,\qquad{\rm at}\ T_{c}\,.$ (3.62) Plugging back into the potential Eq. (3.16) one finds all three terms are the same parametric size $\sim E^{4}T_{c}^{4}/\lambda_{x}^{3}$. At $T_{n}$ which lies just slightly below $T_{c}$, relation Eq. (3.62) approximately holds and thus
$\frac{\phi_{c}}{T_{n}}\approx\frac{2E}{\lambda_{x}}\sim\mathcal{O}(1)\,.$ (3.63)
In a gauge theory one typically has $E\sim g_{x}^{3}$, leading to the following power counting relation for high-temperature phase transitions:
@@
After expanding the one-loop thermal contribution as in Eq. (3.9) and applying the power counting introduced in Eq. (3.64), the leading-order effective potential is obtained as
@@
$(m_{G}^{2}+\Pi_{G})^{3/2}-m_{c}^{3}$ $=\bigl(-\mu_{{\rm eff}}^{2}+\lambda_{x}\phi_{c}^{2}+\xi g_{x}^{2}\phi_{c}^{2}+\tfrac{1}{4}\lambda_{x}T^{2}\bigr)^{3/2}-\bigl(\xi g_{x}^{2}\phi_{c}^{2}\bigr)^{3/2}$
$\approx\bigl(-\mu_{{\rm eff}}^{2}+\lambda_{x}\phi_{c}^{2}+\xi g_{x}^{2}\phi_{c}^{2}\bigr)^{3/2}-\bigl(\xi g_{x}^{2}\phi_{c}^{2}\bigr)^{3/2}\sim\sqrt{\xi}\,\mathcal{O}(g_{x}^{4})\,$, (3.74)
@@
For nucleation-related observables, one typically requires higher accuracy than that provided by the leading order barrier alone. The next-to-leading order (NLO) effective potential is derived within a three-dimensional finite-temperature effective field theory (EFT) framework. The three-dimensional effective potential is computed in a constant background field $\phi_{3}$ using three-dimensional perturbation theory: the leading-order piece is the tree-level potential plus the leading three-dimensional one-loop contributions, while the NLO piece consists of the relevant two-loop vacuum diagrams in the three-dimensional EFT together with the corrections induced by NLO-matched parameters. After renormalization and consistent power counting, one arrives at [47, 48]
@@
The leading-order potential $V_{g_{x}^{3}}^{\rm HT}$, c.f., Eq. (3.69), is manifestly gauge-independent, and the corresponding bounce solution $\phi_{b}$ satisfies
$\nabla^{2}\phi_{b}=\frac{\partial V^{\rm HT}_{g_{x}^{3}}}{\partial\phi_{c}}\biggr|_{\phi_{c}=\phi_{b}}\,,\qquad\phi_{b}(\infty)=0\,,\qquad\phi_{b}^{\prime}(0)=0\,.$ (3.82)
@@
The leading-order effective action is gauge-independent, since the leading-order effective potential $V_{g_{x}^{3}}$ is constructed to be $\xi$-independent.
@@ 3.2.4 Phase transition at low temperatures ($T\ll m_{A^{\prime}}$)
Unlike the zero-temperature toy setup with a positive Higgs mass term, a gauged $U(1)$ Higgs sector contains the standard tachyonic mass-squared term. In the supercooled regime, thermal effects, although being small since $T_{p}\ll v_{x}$, can lift the effective mass term positive and thereby create and maintain a false vacuum at $\phi=0$, as long as the tachyonic mass-squared parameter $\mu_{x}^{2}$ is parametrically small.
@@
Therefore, away from the barrier region, thermal effects are generally negligible since $T\ll\phi_{c}(T)$, and all thermal contributions can be safely neglected relative to the purely field-dependent terms.
Nevertheless, thermal effects are essential for generating the barrier around $\phi\sim T_{p}/g_{x}$, enabling the supercooled phase transition. In the vicinity of $\phi\sim T_{p}/g_{x}$, the temperature-dependent contributions can be comparable to the field-dependent terms. We therefore retain all thermal contributions in both the leading and subleading orders of effective potentials. Consequently, the leading and subleading-order effective potentials are thus written as
@@
$\quad+V_{{\rm daisy}}^{A^{\prime}}(\phi,T)+\frac{T^{4}}{2\pi^{2}}\left[J_{B}\Bigl(\frac{m_{h_{x}}^{2}}{T^{2}}\Bigl)+3J_{B}\Bigl(\frac{m_{A^{\prime}}^{2}}{T^{2}}\Bigl)\right]\,$, (3.96)
@@
$\quad+V_{{\rm daisy}}^{G}(\phi,T)+\frac{T^{4}}{2\pi^{2}}\left[J_{B}\Bigl(\frac{m_{G}^{2}}{T^{2}}\Bigr)-J_{B}\Bigl(\frac{m_{c}^{2}}{T^{2}}\Bigr)\right]\,$, (3.97)
@@
where we incorporate the gauge-independent thermal contributions into the leading-order effective potential, while the $\xi$-dependent thermal terms are assigned to the subleading effective potential. [^6] In the regime $\phi\to 0$ and $T_{p}\gg\phi$, which is most relevant for barrier formation, the high-temperature expansions of the $\xi$-independent terms $V_{{\rm daisy}}^{A^{\prime}}+J_{B}(m_{h_{x}}^{2}/T^{2})+J_{B}(m_{A^{\prime}}^{2}/T^{2})$ appear at lower order in the $g_{x}$ expansion than the $\xi$-dependent terms $V_{{\rm daisy}}^{G}+J_{B}(m_{G}^{2}/T^{2})+J_{B}(m_{c}^{2}/T^{2})$, as discussed in Section 3.2.3. The daisy contribution $V_{{\rm daisy}}^{h_{x}}$ is in general more suppressed than the preceding terms and will therefore be neglected in the following discussion.
@@ 3.2.5 Comments and discussions
  • The power counting rule: For the high-temperature phase transition analysis in [43, 47, 48], the Higgs quartic coupling is taken to scale as $\lambda_{x}\sim g_{x}^{3}$. More precisely, it is assumed to lie in the region $g_{x}^{4}\ll\lambda_{x}\ll g_{x}^{3}$ [3]. [^7] For high-temperature phase transitions, one expects $\frac{g_{x}\phi_{c}}{T_{n}}\sim\frac{g_{x}\phi_{c}}{T_{c}}\ll 1\,.$ (3.107) Using the thermal effective potential in Eq. (3.16), one finds at $T_{c}$, $\frac{\phi_{c}}{T_{c}}\sim\frac{2E}{\lambda_{x}}\sim\frac{g_{x}^{3}}{\lambda_{x}}\,.$ (3.108) Substituting this into the previous condition gives $g_{x}\,\frac{\phi_{c}}{T_{c}}\sim\frac{g_{x}^{4}}{\lambda_{x}}\ll 1\,.$ (3.109) Hence, a consistent high-temperature phase transition typically requires $\lambda_{x}\gg g_{x}^{4}$, also implying that, for fixed $\lambda_{x}$, moving deeper into the high-temperature regime favors smaller values of $g_{x}$. In our parameter scan, we find that all viable high-temperature phase transitions satisfy $\lambda_{x}\lesssim g_{x}^{3}$, while in many cases $\lambda_{x}\ll g_{x}^{4}$. This typically occurs for benchmark points in which $T_{n}$ lies somewhat below $T_{c}$, while the high-temperature approximation $m_{A^{\prime}}/T_{n}\ll 1$ remains reasonably valid.
    Nevertheless, adopting the scaling $\lambda_{x}\sim g_{x}^{3}$ to organize the leading and subleading terms in the potential does yield a gauge-independent effective action, as can be verified analytically using the Nielsen identity. However, alternative scalings such as $\lambda_{x}\sim g_{x}^{4}$ do not lead to consistent results, as the Nielsen identity is not satisfied at the corresponding order, and thus the resulting subleading effective action is not gauge-independent.
@@
  • Gauge parameter dependence: In the gauge-independent treatment, one must ensure that the terms neglected from the full effective potential are indeed parametrically small. Since some of the omitted contributions scale as $\sim\xi^{n}g^{\text{higher powers}}$, the gauge parameter $\xi$ cannot be taken arbitrarily large. Otherwise, the power counting scheme is no longer reliable and the expansion breaks down, even if the Nielsen identity is satisfied for the effective potential up to subleading order.
@@
  • Comments on the low-temperature analysis: Our scan indicates that a supercooled phase transition typically favors a moderately large gauge coupling, e.g., $g_{x}\gtrsim 0.5$, in sharp contrast to the high-temperature case, where smaller $g_{x}$ is generally preferred, c.f, Eq. (3.109). This trend can be understood from two main considerations.
    First, as discussed in the previous section, the low-temperature barrier is generated by the thermal cubic term $\sim T\phi^{3}$ near $\phi=0$, whose coefficient scales as $E\propto g_{x}^{3}$. [^8] The effective thermal potential in Eq. (3.16) is derived in the high-temperature limit, where the $J_{B}$ functions admit a high-temperature expansion. In the low-temperature regime, Eq. (3.16) remains applicable in the field range $\phi\lesssim T_{p}/g_{x}$. In particular, as $\phi\to 0$ the field value is again much smaller than the temperature, and thus the high-temperature expansion remains valid in this small-field region, which is the most relevant for generating a thermal barrier. Nevertheless, Eq. (3.16) and the high-temperature expansions in Eqs. (3.9) and (3.10) should not be applied indiscriminately over the full temperature range in analyses of supercooled phase transitions. A larger $g_{x}$ therefore enhances the cubic term and makes it easier to develop a sufficiently strong barrier for the supercooled phase transition.
    Second, the tree-level tachyonic mass term must be lifted to an effective positive value by thermal mass corrections proportional to $g_{x}^{2}$ in order to create and stabilize a local minimum at $\phi=0$. A larger $g_{x}$ thus keeps the symmetric point $\phi=0$ metastable down to lower temperatures, delaying the loss of metastability, which is essential for realizing a supercooled transition. [^9] A conformal $U(1)$ Higgs sector typically contains no tree-level $\phi^{2}$ term. The local minimum at the origin is instead stabilized by the finite-temperature thermal mass. In this setup, a potential barrier can arise even without relying on the thermal cubic contribution $\sim T\,\phi^{3}$, and may be generated predominantly by the Coleman-Weinberg potential. As a result, the allowed range of the gauge coupling $g_{x}$ can be more flexible than in the conventional gauged $U(1)$ case with a tachyonic mass term. Conformal $U(1)$ models also naturally admit $T_{n,p}\ll v$ and $\alpha\gtrsim 1$, corresponding to strongly supercooled phase transitions that produce enhanced gravitational wave signals with peak frequencies shifted to lower bands. Hence conformal hidden sectors are often discussed as natural pathways to nanohertz band signals.